OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDALS Simplify the number using the imaginary unit I. sqrt( -16

OpenStudy (anonymous):

turn the negative 16 into a positive then square toot....if its a possitive number (ex:5) it would be 5i if not its \[\sqrt{16i}\]

OpenStudy (anonymous):

Another one is Simplify the expression. -5+i ----- 2i

OpenStudy (anonymous):

that one you can just put in you're calculator (-5+i)/(2i)

OpenStudy (anonymous):

Ok so that would be the answer?

OpenStudy (anonymous):

if the awanser is a decimal yes \

OpenStudy (anonymous):

wait how do i do i in my calculator? cuz i think i have to show work

OpenStudy (anonymous):

to show work put it like this... \[\sqrt{-16}\]

OpenStudy (anonymous):

then a arrow going down then put \[\sqrt{16}\]

OpenStudy (anonymous):

simplify it then if its a deciaml leave it as square root of 16 if now put the number then i but with no squareroot

OpenStudy (anonymous):

ok so 4i?

OpenStudy (anonymous):

If that's right then im just confused on the second one

OpenStudy (anonymous):

hello..?

OpenStudy (anonymous):

yes you are correct, and the sceond one you just put into you're calculator as this \[\left( -5+i \right)\div \left( 2i \right)\]